OpenStudy (anonymous):

what is the square route of 81?

OpenStudy (anonymous):

9.

OpenStudy (anonymous):

\[6\sqrt{81}\]

OpenStudy (anonymous):

6*9=54.

OpenStudy (anonymous):

that makes no since?

OpenStudy (anonymous):

do you mean the 6h root of 81?

OpenStudy (anonymous):

\[6\sqrt{81}\]

OpenStudy (anonymous):

6*sqrt(81). sqrt(81)=9. 6*9=54. The 6th root of 81 isn't a whole number.

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

Because 81 is 3^4. 3^4 obviously can't be separated into something to the 6.

OpenStudy (anonymous):

wouldnt it be 3*3*2*2*2*2?

OpenStudy (anonymous):

3*3*2*2*2*2=9*16=144. You're not asking about that...

OpenStudy (anonymous):

ooooohhhhh